Attempt Results Entered Answer Preview Result 26.17 26.17 incorrect 26.27 26.27 incorrect At least one of the answers above is NOT correct 1 point) Sheds R Us sells $900 sheds on a monthly payment plan over 3 years. a) If the interest rate is 3% per month, find the monthly payment. Answer S 26.17 b) If instead the interest rate is 3.25% per month, find the monthly payment Answer $ 26.27
Answer #1

Answer : 3 year = 36 months.

For 36 months the principal amount is $900

For 1 month the principal amount = 900 / 36 = $25

a) Now at interest rate 3% per month;

Interest rate (r) = 3% = 0.03

Time (t) = 3 years = 3/12 months

Principal amount (P) = $900

Simple interest = P * r * t = 900 * 0.03 * 3/12

=> Simple interest = 6.75

Therefore, monthly payment = Monthly principal amount + Simple interest = 25 + 6.75 = $31.75 .

b) At interest rate 3.25% per month;

Interest rate (r) = 3.25% = 3.25 / 100

Simple interest = P * r * t = 900 * 3.25/100 * 3/12

=> Simple interest = 7.31

Therefore, monthly payment = 25 + 7.31 = $32.31 .